# Ngoc BuiY , Shubham SharmaJ, Simran LambaJ, Saumitra MishraJ, Rex YingY

### v1

@@ B.1 Long Generation Evaluation
Experiment settings. For the training, we set the maximum training length to be 16384. We train the retention gates with a learning rate of $2\times 10^{-4}$ and a weight decay of 0.01. For math reasoning tasks, we follow SeerAttn-R [gao2025seerattention] that uses OpenR1-Math-220K [openr1math2025] dataset, which has 564M tokens for training. During training, we use a batch size of 1 for each GPU, and gradient accumulation is set to 4. Other hyperparameters are set to the default in Huggingface Trainer. All training is conducted on 4 H100 GPUs.
Benchmarks. AIME24 [aime2024], GSM8K [cobbe2021training], and MATH-500 [hendrycks2021measuring] are standard math reasoning benchmarks. LongProc [ye2025longproc] is a long-context benchmark of six procedural-generation tasks that require integrating dispersed information and producing structured long-form outputs (up to $\sim 8$K tokens)—from extracting tables from HTML to executing multi-step search to build feasible travel itineraries. The suite spans varied access patterns (sequential vs. targeted retrieval), deductive reasoning demands, and search execution, enabling stress tests of long-range coherence and procedure following. Each task includes deterministic solution procedures and structured outputs, allowing rule-based evaluation (row-level F1 for HTML→TSV, unit tests for pseudocode→code, exact-match traces for Path/ToM, and validators for Countdown/Travel). To probe generation length, we use three difficulty tiers targeting  0.5K, 2K, and 8K output tokens.
@@
Math reasoning results. Figure 6 reports AIME24 performance for Qwen3-1.7B and DeepSeek-R1-Distill variants. Across both families, TRIM-KV consistently outperforms eviction baselines. The gains over heuristic baselines are smaller on DeepSeek-R1-Distill–Llama-8B, which we hypothesize reflects lower attention sparsity in this model compared to the Qwen3 family.
@@
Comparison to a query-agnostic baseline. We provide a comparison to KeyDiff [park2025keydiff], a query-agnostic baseline that only considers a key diversity for eviction. The result in Figure 7 shows that the performance of KeyDiff is significantly worse than that of other baselines. Note that R-KV can be considered as a generalization of KeyDiff since it considers both key diversity and attention scores for eviction heuristics.
@@
Results on LongProc. Table 6 reports KV-eviction results on long procedure–generation tasks. Across tasks and budgets, TRIM-KV achieves the best performance, and it even surpasses the full-cache baseline on Countdown (0.5K/2K) and HTML to TSV (0.5K). Under tighter memory budgets, its margin over heuristic baselines widens.
@@ B.2 Long-Context Evaluation
Experimental settings. We adopt Qwen3-4B-Instruct [qwen3_4b_instruct] as the base model, which supports a context window of up to 256K tokens. Retention gates are trained on a mixture of SynthLong-32K [cerebras2025synthlong], BookSum [kryscinski2021booksum], and Buddhi [singhal2024buddhi], covering sequence lengths from 32K to 128K tokens. We shuffle the combined corpus and train for 10,000 steps (i.e., 10,000 randomly sampled examples), with a maximum training sequence length of 128K and memory capacity $M=4096$. All other settings follow Section 5.1.
Benchmark Dataset. We evaluate chat-assistant capabilities under strict memory budgets using LongMemEvalS [wu2024longmemeval]. This subset provides contexts up to 123k tokens (measured with the Qwen3 tokenizer) and includes six question types that probe long-term memory: single-session-user (SS-User) and single-session-assistant (SS-Assist), which test recall of facts stated by the user or assistant within a session; single-session-preference (SS-Pref), which requires personalized responses from stored personal information; multi-session (Multi), which aggregates or compares information across sessions; knowledge-update (Knowledge), which tracks and answers with the most recent, changed user information; temporal-reasoning (Temporal), which reasons over timestamps and time references.
To evaluate KV-cache eviction methods on this benchmark, we follow the multi-turn, multi-session protocol of [li2024scbench]. Specifically, before each query, the eviction-based model must compress the accumulated dialogue into a fixed-size, reusable KV cache—mirroring real-world assistants that maintain state across turns and sessions under strict memory budgets. We use Qwen3-4B-Instruct [qwen3_4b_instruct] to assess whether model outputs match the ground-truth responses.
Results. The results in Table 7 show that our method outperforms baseline eviction strategies by a significant margin. Especially, TRIM-KV can match the performance of a full cache while using just $25\%$ of the KV budget.
@@ B.3 Chunked-Prefill Evaluation
In this section, we evaluate our method in the chunked-prefill setting [huang2024locret], which enables long-context inference without exceeding memory limits. In this framework, long prompts are split into multiple chunks; the model computes the KV cache for each chunk sequentially and compresses the cache whenever it surpasses the memory budget. We compare our method against LocRet [huang2024locret], a learnable KV eviction baseline that also assigns token-importance scores for eviction. Following the experimental setup of LocRet, we evaluate TRIM-KV on the LongBench [bai2024longbench] and LongBench-V2 [bai2025longbench] benchmarks. For LongBench-V2, we restrict evaluation to examples with context length below 128K tokens, matching the maximum context length advertised for Phi3-mini-128K. To train TRIM-KV, we use only LongAlpaca [chen2023longlora], as in LocRet, to ensure that improvements are not attributable to differences in training data. In this setting, we set $M=2048$, and keep all other hyperparameters identical to those in Section 5.2.
For a fair comparison, we adopt the hyperparameters used by LocRet: the chunk size is set to 3072 and the budget size to 6000. We evaluate performance on Phi-3-mini-128K [abdin2024phi], reproducing Table 6 from the LocRet paper. Since the original LongBench experiments have not been released, we use the default chat template of Phi-3-mini-128K for all runs. Table 8 and 9 report the results for LongBench [bai2024longbench] and LongBench-V2 [bai2025longbench] benchmarks, respectively. We observe discrepancies in the full-KV performance, likely due to differences in chat templates.
@@
Table 8: Performance on long-context tasks with Phi3-mini-128K on the LongBench benchmark, including average relative change (Avg. $\Delta$) compared to Full KV. * indicates that the numbers are reported from [huang2024locret, Table 6].
Method Acc. Short Acc. Medium Acc. Easy Acc. Hard Avg. Acc $\Delta$ (%)
@@ B.4 Additional Ablation Studies
Ablation on training datasets. In Section 5.1, we train the retention gates on a reasoning dataset—OpenR1-Math [openr1math2025]—and evaluate on AIME24, MATH-500, and GSM8K. This follows standard practice and matches the setting of [gao2025seerattention], ensuring a fair comparison. To assess cross-domain generalization, we instead train the gates on general long-context datasets (SynthLong, BookSum, Buddhi), similar to Section 5.2, and then evaluate on math reasoning benchmarks to test whether retention scores learned from general data transfer to long chain-of-thought tasks. As shown in Figure 8, gates trained on general datasets generalize well and even surpass math-specific training at a 2048 KV budget. However, their performance degrades more quickly under tighter KV budgets. Overall, these results are promising and suggest that scaling the training of the retention gates by combining all datasets could yield further improvements.
Figure 9: Ablating different objective components.
Ablation for the objective function. We ablate the objective by training the Qwen3-4B retention gates with different loss combinations and report AIME24 pass@1 at a 4096-token KV budget in Figure 9. Here, we consider both forward KL divergence and reversed KL divergence for distillation loss. Generally, all distillation losses perform well on their own. However, reversed KL underperforms when compared to forward KL. Both show benefits in combination with next token prediction. The memory capacity loss is essential for compression, and removing it leads to a sharp drop.
@@ Appendix C Additional Qualitative Results
Figure 14: Visualization of token retention scores $\beta_{i}$ from the retention gate in layer 15, head 2 of the Qwen3-4B model. Brighter colors denote higher retention; gray indicates near-zero retention. Observation: This head consistently assigns a high retention to the problem description and tokens expressing mathematical operations, suggesting a specialization for task-critical content. In contrast, filler phrases (e.g., ‘wait’, ‘let me think’) receive very low retention. Moreover, for a multi-digit number, retaining only the hidden state of the final digit suffices to maintain performance, suggesting that the last digit’s representation already captures the semantic information of the whole number.
Figure 15: Visualization of token retention scores $\beta_{i}$ from the retention gate in layer 9, head 7 of the Qwen3-4B model. Brighter colors denote higher retention; gray indicates near-zero retention. Observations: This head exhibits a retention pattern similar to layer 15, head 2, but places greater emphasis on arithmetic operators (e.g., +,,/+,-,/).
Figure 16: Visualization of token retention scores $\beta_{i}$ from the retention gate in layer 35, head 2 of the Qwen3-4B model. Brighter colors denote higher retention; gray indicates near-zero retention. Observation: Unlike layer 15, head 2 and layer 9, head 7, this head assigns elevated retention to general-purpose tokens that support coherence and instruction following; for example, commands and the directive boxed{) receive high scores, while tokens associated with mathematical operations receive low retention.
Figure 17: Visualization of tokens retained in the layer 0 head 3 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 18: Visualization of tokens retained in the layer 0 head 6 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
@@
Figure 20: Visualization of tokens retained in the layer 15 head 2 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 21: Visualization of tokens retained in the layer 16 head 6 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 22: Visualization of tokens retained in the layer 30 head 4 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens. Observations: This head mostly keeps period tokens. This suggests that this head may implicitly perform gist tokens that summarize information from the previous sentence. This contrasts with recent trends that advocate for saving a chunk of tokens [yuan2025native, gao2025seerattention]. Our results suggest that it can be more budget-efficient to save individual tokens because they already capture contextual information.
Figure 23: Visualization of tokens retained in the layer 34 head 1 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 24: Visualization of tokens retained in the layer 35 head 3 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 25: Visualization of tokens retained in the layer 35 head 5 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 26: Visualization of tokens retained in the layer 35 head 7 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.

### v2

@@ B.1 Long Generation Evaluation
Experiment settings. For the training, we set the maximum training length to be 16384. We train the retention gates with a learning rate of $2\times 10^{-4}$ and a weight decay of 0.01. For math reasoning tasks, we follow SeerAttn-R [Gao et al., 2025] that uses OpenR1-Math-220K [Hugging Face, 2025] dataset, which has 564M tokens for training. During training, we use a batch size of 1 for each GPU, and gradient accumulation is set to 4. Other hyperparameters are set to the default in Huggingface Trainer. All training is conducted on 4 H100 GPUs.
Benchmarks. AIME24 [Art of Problem Solving, 2024], GSM8K [Cobbe et al., 2021], and MATH-500 [Hendrycks et al., 2021] are standard math reasoning benchmarks. LongProc [Ye et al., 2025] is a long-context benchmark of six procedural-generation tasks that require integrating dispersed information and producing structured long-form outputs (up to $\sim 8$K tokens)—from extracting tables from HTML to executing multi-step search to build feasible travel itineraries. The suite spans varied access patterns (sequential vs. targeted retrieval), deductive reasoning demands, and search execution, enabling stress tests of long-range coherence and procedure following. Each task includes deterministic solution procedures and structured outputs, allowing rule-based evaluation (row-level F1 for HTML→TSV, unit tests for pseudocode→code, exact-match traces for Path/ToM, and validators for Countdown/Travel). To probe generation length, we use three difficulty tiers targeting  0.5K, 2K, and 8K output tokens.
@@
Math reasoning results. Figure 6 reports AIME24 performance for Qwen3-1.7B and DeepSeek-R1-Distill variants. Across both families, TRIM-KV consistently outperforms eviction baselines.
@@
Comparison to a query-agnostic baseline. We provide a comparison to KeyDiff [Park et al., 2025], a query-agnostic baseline that only considers a key diversity for eviction. The result in Figure 7 shows that the performance of KeyDiff is significantly worse than that of other baselines. Note that R-KV can be considered as a generalization of KeyDiff since it considers both key diversity and attention scores for eviction heuristics.
@@
Results on LongProc. Table 7 reports KV-eviction results on long procedure–generation tasks. Across tasks and budgets, TRIM-KV achieves the best performance, and it even surpasses the full-cache baseline on Countdown (0.5K/2K). Under tighter memory budgets, its margin over heuristic baselines widens.
@@ B.2 Long-Context Evaluation
Experimental settings. We adopt Qwen3-4B-Instruct [Qwen3, 2025] as the base model, which supports a context window of up to 256K tokens. Retention gates are trained on a mixture of SynthLong-32K [Lazarevich et al., 2025], BookSum [Kryściński et al., 2021], and Buddhi [Singhal, 2024], covering sequence lengths from 32K to 128K tokens. We shuffle the combined corpus and train for 10,000 steps (i.e., 10,000 randomly sampled examples), with a maximum training sequence length of 128K and memory capacity $M=1024$. All other settings follow Section 5.1.
Benchmark Dataset. We evaluate chat-assistant capabilities under strict memory budgets using LongMemEvalS [Wu et al., 2024a]. This subset provides contexts up to 123k tokens (measured with the Qwen3 tokenizer) and includes six question types that probe long-term memory: single-session-user (SS-User) and single-session-assistant (SS-Assist), which test recall of facts stated by the user or assistant within a session; single-session-preference (SS-Pref), which requires personalized responses from stored personal information; multi-session (Multi), which aggregates or compares information across sessions; knowledge-update (Knowledge), which tracks and answers with the most recent, changed user information; temporal-reasoning (Temporal), which reasons over timestamps and time references.
To evaluate KV-cache eviction methods on this benchmark, we follow the multi-turn, multi-session protocol of [Li et al., 2024b]. Specifically, before each query, the eviction-based model must compress the accumulated dialogue into a fixed-size, reusable KV cache—mirroring real-world assistants that maintain state across turns and sessions under strict memory budgets. We use Qwen3-4B-Instruct [Qwen3, 2025] to assess whether model outputs match the ground-truth responses.
Results. The results in Table 8 show that our method outperforms baseline eviction strategies by a significant margin.
@@ B.3 Chunked-Prefill Evaluation
In this section, we evaluate our method in the chunked-prefill setting [Huang et al., 2024], which enables long-context inference without exceeding memory limits. In this framework, long prompts are split into multiple chunks; the model computes the KV cache for each chunk sequentially and compresses the cache whenever it surpasses the memory budget. We compare our method against LocRet [Huang et al., 2024], a learnable KV eviction baseline that also assigns token-importance scores for eviction. Following the experimental setup of LocRet, we evaluate TRIM-KV on the LongBench [Bai et al., 2024] and LongBench-V2 [Bai et al., 2025] benchmarks. For LongBench-V2, we restrict evaluation to examples with context length below 128K tokens, matching the maximum context length advertised for Phi3-mini-128K. To train TRIM-KV, we use only LongAlpaca [Chen et al., 2023], as in LocRet, to ensure that improvements are not attributable to differences in training data. In this setting, we set $M=512$, and keep all other hyperparameters identical to those in Section 5.2.
For a fair comparison, we adopt the hyperparameters used by LocRet: the chunk size is set to 3072 and the budget size to 6000. We evaluate performance on Phi-3-mini-128K [Abdin et al., 2024], reproducing Table 6 from the LocRet paper. Since the original LongBench experiments have not been released, we use the default chat template of Phi-3-mini-128K for all runs. Table 9 and 10 report the results for LongBench [Bai et al., 2024] and LongBench-V2 [Bai et al., 2025] benchmarks, respectively. We observe discrepancies in the full-KV performance, likely due to differences in chat templates.
@@
Table 9: Performance on long-context tasks with Phi3-mini-128K on the LongBench benchmark, including average relative change (Avg. $\Delta$) compared to Full KV. * indicates that the numbers are reported from [Huang et al., 2024, Table 6].
Method Acc. Short Acc. Medium Acc. Easy Acc. Hard Avg. Acc Avg. $\Delta$ (%)
@@ B.4 Additional Ablation Studies
Ablation on training datasets. In Section 5.1, we train the retention gates on a reasoning dataset—OpenR1-Math [Hugging Face, 2025]—and evaluate on AIME24, MATH-500, and GSM8K. This follows standard practice and matches the setting of [Gao et al., 2025], ensuring a fair comparison. To assess cross-domain generalization, we instead train the gates on general long-context datasets (SynthLong, BookSum, Buddhi), similar to Section 5.2, and then evaluate on math reasoning benchmarks to test whether retention scores learned from general data transfer to long chain-of-thought tasks. As shown in Figure 8, gates trained on general datasets generalize well and even surpass math-specific training at a 2048 KV budget. However, their performance degrades more quickly under tighter KV budgets. Overall, these results are promising and suggest that scaling the training of the retention gates by combining all datasets could yield further improvements.
@@ Appendix C Additional Qualitative Results
Figure 13: Visualization of tokens retained in the layer 4 head 5 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 14: Visualization of tokens retained in the layer 6 head 4 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 15: Visualization of tokens retained in the layer 7 head 3 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
@@
Figure 17: Visualization of tokens retained in the layer 15 head 5 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.
Figure 18: Visualization of tokens retained in the layer 27 head 2 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens. Observations: This head mostly keeps period tokens. This suggests that this head may implicitly perform gist tokens that summarize information from the previous sentence. This contrasts with recent trends that advocate for saving a chunk of tokens [Yuan et al., 2025, Gao et al., 2025]. Our results suggest that it can be more budget-efficient to save individual tokens because they already capture contextual information.
Figure 19: Visualization of tokens retained in the layer 32 head 2 of the KV cache after generation, where the KV budget is 256. The model is Qwen3-4B. Bold blue indicates tokens retained in the KV cache, where gray indicates evicted tokens.